Find the pH of a 0.100 M solution of a weak monoprotic acid having Ka= 1.2×10−3.

Express your answer to two decimal places.
Answer #1

Let the acid be HA

HA----> H+ +A-

Initial   HA = 0.1    H+ = [A-]= 0

change [HA] =-x   [H+] =[A-] =x

At Equilibrium   HA= 1-x       [H+] =[A-] =x

Ka= x2/(1-x)= 1.2*10-3

this equation gives when solved usong solved gives x =0.0341

[H+] =0.0341

pH= -log(0.0341)=1.467